# Disentangling Boltzmann brains, the time-asymmetry of memory, the H theorem, and the second law

### v2

@@ Abstract
Are you, with your perceptions, memories and observational data, a Boltzmann brain, i.e., a statistical fluctuation out of the thermal equilibrium of the universe? Arguments are given in the literature for and against taking this hypothesis seriously. Complicating these analyses have been the many subtle – and very often implicit – entanglements between related arguments that have been given for the past hypothesis, the second law, and even Bayesian inference of the reliability of experimental data. These entanglements can easily lead to circular reasoning. To help disentangle those arguments, since almost all of them involve Boltzmann’s H theorem, we begin by formalizing the H theorem as a time-symmetric, time-translation invariant Markov process over the entropy values of the universe. Crucially, this process does not specify the time(s) on which we should condition it in order to infer the stochastic dynamics of our universe’s entropy. Any such choice of conditioning events must be introduced as an independent assumption. This observation allows us to disentangle the standard Boltzmann brain hypothesis, its “1000CE” variant, the past hypothesis, the second law, and the reliability of our experimental data, all in a fully formal manner. In particular, we show that they all adopt the $H$ theorem’s stipulation that the universe’s entropy evolves as a Markov processes, and all make an arbitrary assumption that the process should be conditioned on a single moment in time. Their only difference is what single time to condition on. In this aspect, the Boltzmann brain hypothesis and the second law are equally legitimate (or not).
@@ I Introduction
“Boltzmann brain” (BB) is the name given to a phenomenon that is in principle allowed in statistical mechanics. Consider a large statistical system formed by a mixture of particles of different kinds that remains in thermal equilibrium for an arbitrarily long span of time. According to statistical mechanics, there are fluctuations at thermal equilibrium, and in principle all configurations can be reached by such fluctuations with enough time available. Consider one of these random fluctuations giving rise – just by chance – precisely to a brain like ours, complete with our memories, information and perception, fluctuating for a brief moment into existence. That would be a BB (see for instance Albrecht and Sorbo (2004); Linde (2007); Aguirre et al. (2012)). Such a brain would have exactly our worldviews, memories, perceptions, and it would know and feel precisely what we know and the way we feel right now. Is it possible to know that we are not a BB?
An intuitive answer is that we are not, since the probability for such a fluctuation to occur is fantastically small, and the expected time for this to happen is colossally longer than the current age of the universe. However, recent literature Albrecht and Sorbo (2004); Linde (2007); Goheer et al. (2003); Dyson et al. (2002); Page (2008) has questioned this simple answer, presenting an argument that current science makes the possibility of being a BB harder to dispel. Here, we introduce a weak point of the BB argument (for a similar analysis, see Myrvold (2016)).
The argument informally holds that the BB conclusion follows from two pieces of knowledge:
@@
    The set $\cal L$ of all the mechanical laws of physics that we know, which are symmetric under time inversion.
@@
    The set $\cal D$ of all present observations (or Data) that we have, which includes all records, recorded data, and memories existing in the present.
What can then we say about the world, based on our data $\cal D$? The set of laws $\cal L$ admits a large family of solutions. Let us restrict ourselves to these solutions to those compatible with $\cal D$, in the sense that $\cal D$ has non-infinitesimal likelihood under those solutions. Ignoring the other solutions, for simplicity, let us assign equal probability to all these solutions that we restrict ourselves to.
@@
The dynamical laws (with the equiprobability assumption) imply the Boltzmann’s H theorem, which says that entropy increases both into the past and into the future from any given time at which entropy is known to have some specific low value. This theorem is time-symmetric about this special time. The standard formulation of the second law of thermodynamics says that we are in the future of such a time. Note though that the most probable situation, given the current observations, is that we happen to be at that special point, not in its future. In other words, the most probable situation is that we are just an entropy fluctuation, which is to say that we are a BB.
One may respond to this argument that we also know the second law of thermodynamics, and from this we know that entropy grows with $t$, i.e., that the special point with low given entropy was in our past. But our knowledge of the second law arose from consideration of our data records about the past: how do we know that these are data records of the past and aren’t themselves due to a fluctuation?
We commonly assume that entropy was in fact low in the past and has been growing since. But assuming that the entropy in the past was (sufficiently) low amounts to a very improbable assumption, with respect to the equiprobability of the solutions compatible with $\cal D$. On the basis of this equiprobability, the data $\cal D$ are more likely to be the result of a fluctuation than to be the result of low entropy in the distant past. This is simple to show: the entropy today is higher than this assumed past entropy and entropy is directly related to probability in this sense. Low past entropy, in this logic, is not an account of why the present is out of equilibrium: it is additional in-probability added to how special the present is.
Additionally, assuming ergodicity, if the set of laws $\cal L$ admits thermalization in the distant future and the universe settles into a long living equilibrium state (Boltzmann’s “thermal death”), then even if the universe started off in a low entropy state, still $\cal D$ can be realized not once but an infinite number of times (up to arbitrary precision) in the distant future, making it more reasonable to infer that an observation of $\cal D$ is one of these fluctuations, rather than something realized in the initial thermalization transient. These arguments appear to converge on the conclusion that we should take the hypothesis of being BB seriously.
Yet, there is something strident and unconvincing in all this. Is this just a resistance due to our wrong naive intuition, or is there a real problem in the argument above? We think that the second case is the correct one, and here we point out where one difficulty is.
First, we will explain that the standard argument in favor of the BB hypothesis contradicts itself by implicitly assuming the second law, which is our first contribution. This is semi-formal only, and we are not overly scrupulous of the precise Bayesian foundations of the philosophy of science. A more detailed, formal, and general version of our argument is presented later in the paper. Our goal is to highlight a flaw in the standard argument, while clarifying why that flaw doesn’t apply to our current understanding of the laws of the universe.
Before showing this argument, there is a preliminary ambiguity to dispel. It is different to argue that something is unlikely, even extremely unlikely, from arguing that it is impossible. Nothing is impossible unless it violates logic. This idea was well-summarized long ago by Zhuang Zi Ziporyn et al. (2020) in the famous apologue of the dream of the butterfly: Zhuang Zi dreamt of being a butterfly happily flying around among flowers. When he woke up, he puzzled: “How can I know if I am Zhuang Zi who dreamt of being a butterfly, or I am a butterfly dreaming of being Zhuang Zi?”.
But all of these possibilities, precisely because they are innumerable and of any possible sort, are typically dismissed in science. The standard view is that they form a grey ocean of nonsense, all equally irrelevant, unless something makes one of them plausible. There is no totally rigorous argument to exclude any of them. In fact, no data that we could acquire could confirm that Zhuang Zi is a butterfly dreaming — nor could any data refute that hypothesis, establishing that he is not just such a butterfly. This is why such a hypothesis is consigned to the “grey ocean”, under the standard view of science. In this same, trivial sense, being a BB cannot be rigorously excluded, but rather would be consigned to the “grey ocean” under the standard view of science. The puzzle of BBs is not about the abstract possibility of being a fluctuation: it is about the probability of being a fluctuation. It is not about what is logically possible: it is about what we say about how probable it is. What the standard argument is supposed to argue for is that the probability of the BB hypothesis is close to 1.
@@
This standard argument for the BB hypothesis initially arose in the context of de Sitter quantum cosmology Albrecht and Sorbo (2004); Linde (2007); Goheer et al. (2003); Dyson et al. (2002). If such a universe evolves into the future, it asymptotes to a vacuum state with a finite, non-zero temperature. This equilibrium state has fluctuations that act similarly to thermal fluctuations in a classical gas at equilibrium, which satisfy an H theorem and corresponding recurrence theorems. Many have maintained that the laws of physics must be able to suppress this possibility of a BB. There have been plenty of criticisms and arguments against the standard BB argument Norton (2015); Carroll (2020); Elga (forthcoming); Wallace (2023a); Page (2024); Dogramaci and Schoenfield (2025). Many fall into categories related to instability Carroll (2020); Elga (forthcoming), Bayesian analysis Wallace (2023a); Page (2024); Dogramaci and Schoenfield (2025) and physics-dependent “loopholes” arising from the precise nature of cosmology and quantum mechanics (see Carroll (2020) and references therein).
@@
In this paper, we do not dispute the major role of cosmology in the second law of thermodynamics Rovelli (2019); Davies (1977); Scharnhorst and Aguirre (2023) or investigate precisely which models predict BB and which do not. Indeed, there are even major arguments that a framework for studying BB depending strictly on the H theorem, without cosmology, is incomplete. It has long been known that the BB issue is highly sensitive to schema for assigning probabilities to values of the scale factor in FLRW scalar field cosmology models Page (2008); De Simone et al. (2010); Vilenkin (2007). We simply start from the working basis of pure statistical physics to unify and contextualize a recent genre of arguments in the literature.
Despite the wide attention BBs have received in the literature and the implicit role of memory in those arguments, analyses have yet to incorporate a precise treatment of memory in physical systems.
In fact, the notion of physical instantiation of memory is central to our arguments in this paper, and we believe to understanding many BB criticisms. (By criticism, we refer to any argument that concludes that the probability of the BB hypothesis is at or near zero). In Reichenbach (1991); Rovelli (2024), Reichenbach elucidates a prototype for memory in physical systems: branch systems. These are a type of physical system with a second law and many weakly interacting subsystems, each with its own thermalization rate but all “branching off” from a low-entropy system at an earlier time. A branch system can produce plentiful records. Recently, physical memory has been formalized in Wolpert (1992); Rovelli (2022); Wolpert and Kipper (2024). Reference Wolpert and Kipper (2024) defines a memory system as any physical system $M$ whose present state $(M_{0},t_{0})$ has information about a different systems $W$ state $(W_{1},t_{1})$ at a previous time. In order for the system $M$ to qualify as being memory for system $W$, there must be a sufficiently high amount of mutual information between the two states $(M_{0},t_{0})$ and $(W_{1},t_{1})$. Critically, Wolpert (1992); Rovelli (2022); Wolpert and Kipper (2024) all emphasize that some notions of memory, specifically Type-3 of Wolpert and Kipper (2024), rely on the second law itself. Type-3 systems rely on the second law to return the system to its “ready state.” Examples of such a state are a smooth beach after footprints wash away. In practice, all records of past physical experiments rely on Type-3 memory systems. In turn, note that any explanation of the second law invoking temporally asymmetric memory (e.g. Type-3) is fundamentally circular. Aside from these key insights, we do not concern ourselves with the formal details of memory systems in this paper. Rather, we focus on the reliability of properly defined experimental data, with the understanding that reliability itself requires the second law.
The argument in the next section is meant to intuitively convey the necessity of discussing the reliability, which we present in a semi-formal manner. However, some might object that this argument only establishes that the priors humans use to reason about the past are inconsistent, and that such an argument does not have any implications for whether or not the implications of those priors are valid (e.g., due to other reasons). Others might object that what we present there is a slight variant of other informal arguments that have been made in the literature. To address these concerns, in Section III we present our argument in a fully formal manner. In particular, we precisely specify the nature of the data $\cal{D}$ (also see Section II) that grounds the BB hypothesis. As elaborated above, $\cal{D}$ generally contains data that implies the laws of physics and separate data that records current or previous entropy values.
We suggest an alternate to the BB hypothesis, called the 1000CE hypothesis, which shows that the domain of BB-type problems is larger than the standard argument suggests. The framework is this paper is powerful enough to investigate the domain of these problems. In particular, it is far more rigorous about the Bayesian reasoning used when we reason about the past than any earlier work in the literature. As we show, a fully formal analysis relies centrally on some subtleties concerning the (mis)application of Boltzmann’s H theorem to reason about the past, which were first investigated in Scharnhorst et al. (2024). We end with a discussion.
II The flaw in the standard arguments for Boltzmann Brains
Recall the standard argument for the BB hypothesis: given knowledge of i) the laws $\cal L$ and ii) the present data of observation $\cal D$, it follows that the BB hypothesis ($\cal B$) has high probability. Let us formalize this result by writing that the probability $P({\cal B}\,|\,{\cal L},{\cal D})$ of the BB hypothesis, given the laws of physics and the data observed in the present, is near unit. We can also write this as
@@ II The flaw in the standard arguments for Boltzmann Brains
Indeed, the standard argument goes through when properly formalized (see Section III). The problem we intend to point out is not in the logic of the argument: it is in the premises $\cal L$ and $\cal D$. And it is not either than (i) and (ii) are individually wrong. It is that we infer $\cal L$ from $\cal D$ — but that inference in turn invokes those very laws $\cal L$ we wish to infer. (Specifically, it relies on our using the second law of thermodynamics Wolpert and Kipper (2024); Wolpert (1992).) So we have circular reasoning. Viewed differently, Both (i) and (ii) are right. And yet, the argument is incorrect. The problem can be seen as arising because (i) and (ii) are incomplete, and likelihood derived from incomplete premises can be strongly misleading.
To show that this is the case, let us first make a general observation. From a certain amount of information $I$ that we judge reliable, we can deduce some consequences. If we restrict to a subset $i\subset I$ of the information, we may deduce some consequences that are clearly false in the light of $I$. For instance, suppose that the information $i$ is a video of a closed camera circuit, from which we see that John was killed at 6pm in his house, and at 6pm Bob, upset because John had kissed his girlfriend, was in the same house, and forensic evidence shows that the gun, with Bob fingerprints, is the weapon that killed John. This seems to be strong evidence against Bob. But imagine that there is also a second video, where we see that at 6pm a robber had entered John’s house, taken the gun from Bob’s coat hanging in the entrance, and shot John while he was discussing in a civilized manner the jealously issue with his dear friend Bob. Then the evidence on the basis of $i$ appears to be misleading, in the light of the larger evidence $I$. A shadowy secret service could have Bob convicted by carefully hiding the second video recorded by the closed circuit camera. All evidence can be doubted, but doubting selectively leads to shaky conclusions.
Let us put this in a bit more formal terms. The likelihood that Bob is culpable ($B$), given the evidence provided by the two tapes $T_{1}$ and $T_{2}$ together is negative:
$(T_{1},T_{2})\to({\rm not}\ B)$ (2)
And yet, the likelihood that that Bob is culpable given the evidence provided by the incomplete evidence $T_{1}$ alone is high
$T_{1}\to B$ (3)
(and misleading). Succinctly, we can write the probability of a proposition $X$, given a condition $Y$ that highly suggests $X$ as $P(X\,|\,Y)\lesssim 1$ but $P(X\,|\,Y,Z)=0$, for a suitable condition $Z$. $Z$ could be, but is not limited, to $\neg X$. There is no contradiction. Incomplete evidence can suggest high probability to false conclusions.
We are now going to argue that the two conditions (i) and (ii) are incomplete evidence, and the BB hypothesis is like Bob’s culpability: a false conclusion from selective incomplete evidence. To this aim, let’s focus on the assumption (i): the laws $\cal L$. How do we know the laws of the universe that we claim to know? By knowing only some present data $\cal D$, i.e., properties of the current universe, we cannot infer these laws. To make such an inference, we must know some data concerning the properties of the universe in the past, $\cal P$. This is because the laws $\cal L$ are relations between the values of the physical variables at different times, hence we cannot know or infer anything about them unless we also know data about the world at a time different that the present. That is, we have
@@
How do we access the past data $\cal P$? As described in Wolpert and Kipper (2024), to do this we have to trust the reliability of records or memories we have at the present that concern these past data. Let us call $\cal R$ the assumption that such memories of past data are statistically reliable. [^1] In the terminology of Wolpert and Kipper (2024), the results of almost all science experiments are recorded in a “type-3” memory. The second law can be applied to deduce information about the past state of the universe from the current state of such a memory, i.e., such memories are “reliable”. Accordingly, we can formalize $\cal R$ as the statement that we have a type-3 memory, whose current state is specified in $\cal D$, and that provides the information $\cal P$ about the past state of the universe. For notational simplicity, we indicate these memories as parts of $\cal D$, so that the combination $(\cal{D},\cal{R})$ establishes $\cal P$. So whenever we write $(\cal{D},\cal{R})$, that is equivalent to writing $(\cal{D},\cal{R},\cal{P})$.
Now, let us consider the two cases separately: whether we make the assumption $\cal R$ or we do not.
@@
which is the opposite of what is given by the incomplete evidence. Rewriting in terms of probabilities, (6) means that $P({\cal L}\,|\,{\cal{D}},{\cal R})$ is close to $1$. The chain rule for probability says
$P({\cal B},\cal{L}\,|\,{\cal{D}},{\cal R})=P({\cal B}\,|\,{\cal{D}},{\cal R},{\cal L})P({\cal L}\,|\,{\cal{D}},{\cal R}),$ (7)
which is $\simeq 0$ since $P({\cal B}\,|\,{\cal{D}},{\cal R},{\cal L})\simeq 0$. Combining this with the fact that $P({\cal B},\cal{L}\,|\,{\cal{D}},{\cal R})\sim P({\cal B}\,|\,{\cal{D}},{\cal R})$, we have $P({\cal B}\,|\,{\cal{D}},{\cal R})\simeq 0$.
Therefore, returning to our simplified notation,
@@
In order to have (reliable) memories or traces we need dissipation towards the future, namely increasing entropy in the time before the present Wolpert and Kipper (2024); Rovelli (2022), while the BB hypothesis assumes that entropy decreases in this time (to lead to the large fluctuation). That is, assuming that traces are reliable yields no BB. This is precisely the case of incomplete evidence. The conclusion of the standard argument for BBs is false, given reliability: it is an illusion due to selected incomplete evidence. Taking account all available evidence voids it.
What about if we instead do not assume $\cal R$? In this case, we have no argument in support of $\cal L$. Of course $\cal B$ does not follow from $\cal D$ alone, without $\cal L$. Without an assumption about reliability of memories, we have no reasons to believe in BBs. Data at a single moment of time are compatible with any dynamical law, namely with both $\cal{B}$ and $\neg\cal{B}$. Absent other assumptions, they characterize the state of the universe at a single time, at best. They are not informative about the laws of the universe.
One may object: we can consider $\cal L$ not as a component of our knowledge, but rather as a fact of the world. That is, suppose we say: the world is truly governed by the laws $\cal L$, irrespectively from our knowledge. What can then we deduce from this fact conjoint with the fact that we know $\cal D$? This appears to circumvent the observation that our knowledge of $\cal L$ depends on our knowledge of records and on the assumption $\cal R$. But this objection does not hold. Assigning likelihood to an event is a subjective matter, not an objective matter that pertains solely to the event. By themselves, events either happen or don’t. Facts are either true of false. To reason about likelihood can only be based on incomplete knowledge and assumptions. The only reasonable question here is the likelihood that we assign to the possibility of being of BBs, and this only depends on what we consider reliable. Our actual considering $\cal L$ reliable is grounded into our considering records reliable, hence not being in a fluctuation. That is, necessitates the assumption $\cal R$.
This discussion may leave a sense of confusion. How do we actually know the laws? Don’t we derive them from current observations? The answer is yes, but not directly. In the standard practice of science, the answer is evident: we do rely on data at different times. We do assume memories and traces to be at least to some extent reliable. When we are at a given moment of time, how can we access data at different times? For this we need them to be recorded and their records to be all present to us in a reliable manner. For this, we have to assume that the records and our memory are reliable. This is all to say that justifying a prior for $\cal{L}$ without reliable data seems very difficult, but if one justifies such a prior without assuming reliability, the BB hypothesis can be rescued in some sense (see Section III.3).
One might characterize these considerations as moving the BB hypothesis back into the vast realm of plausibility - not likelihood. It is possible that the BB hypothesis is true and this text does not exist and is instantaneously hallucinated by neurons randomly happening to fluctuate into this perception. But in many respects, this hypothesis is as likely as being the dream of a butterfly.
So, we have concluded that the reliability of experimental data is central to analysis of the BB hypothesis (in addition to specifying precisely what that data is). How can we go from this together with the H theorem, formally and carefully in a fully Bayesian manner, to fully disentangle the relationship between the second law, BBs, and how it is that we have memory in first place? As we show in the rest of this paper, Scharnhorst et al. (2024) provides a crucial framework to properly understand the relationships between these concepts.
III Relation between the H theorem, Boltzmann Brains, and the second law
The discussion in the preceding sections was intended to convey our central argument in an intuitive way. Since that argument seems to involve circular reasoning, some might object that it only establishes that our priors are inconsistent (see Appendix B). This would mean that that argument don’t tell us anything about how valid its conclusions are, just that the priors it relies on don’t themselves establish those conclusions. Others might object that what we presented above is only a slight variant of other informal arguments that have been made before, e.g. Carroll (2020).

### v3

# Disentangling Boltzmann brains, the time-asymmetry of memory, and the second law

@@ Abstract
Are your perceptions, memories and observational data, a statistical fluctuation out of the thermal equilibrium of the universe, having no correlation with the actual past state of the universe? Arguments are given in the literature for and against this “Boltzmann brain” hypothesis. Complicating these arguments have been the many subtle – and very often implicit – joint dependencies among these arguments and others that have been given for the past hypothesis, the second law, and even for Bayesian inference of the reliability of experimental data. These dependencies can easily lead to circular reasoning. To avoid this problem, since all of these arguments involve the stochastic properties of the dynamics of the universe’s entropy, we begin by formalizing that dynamics as a time-symmetric, time-translation invariant Markov process, which we call the entropy conjecture. Crucially, like all stochastic processes, the entropy conjecture does not specify any time(s) which it should be conditioned on in order to infer the stochastic dynamics of our universe’s entropy. Any such choice of conditioning times and associated entropy values must be introduced as an independent assumption. This observation allows us to disentangle the standard Boltzmann brain hypothesis, its “1000CE” variant, the past hypothesis, the second law, and the reliability of our experimental data, all in a fully formal manner. In particular, we show that these all make an arbitrary assumption that the dynamics of the universe’s entropy should be conditioned on a single event at a single moment in time, differing only in the details of their assumptions. In this aspect, the Boltzmann brain hypothesis and the second law are equally legitimate (or not).
@@ I Introduction
Consider a large statistical system formed by a mixture of particles of different kinds that remains in thermal equilibrium for an arbitrarily long span of time. According to statistical mechanics, there are fluctuations at thermal equilibrium, and in principle all configurations can be reached by such fluctuations with enough time available. Consider one of these random fluctuations giving rise – just by chance – precisely to a brain like ours, complete with our memories, information and perception, but no corresponding external physical world. Such a brain – indistinguishable from ours, but arising as a fleeting fluctuation – is call a “Boltzmann brain” (BB) Albrecht and Sorbo (2004); Linde (2007); Aguirre et al. (2012)). Such a brain would have exactly our current memories, perceptions, and it would know and feel precisely what we know and feel right now.
Is it possible to know that we are not a BB? An intuitive answer is that we yes, we can know this, that we are not a BB, since the probability for such a fluctuation to occur is fantastically small, and the expected time for this to happen is colossally longer than the current age of the universe. However, recent literature Albrecht and Sorbo (2004); Linde (2007); Goheer et al. (2003); Dyson et al. (2002); Page (2008) has questioned this simple answer. Counter-counter arguments have then said that the reasoning behind the BB hypothesis is “cognitively unstable”, that in fact it is “unstable reasoning”Myrvold (2016); Carroll (2016); Elga (2025); Wallace (2023a).
To understand the issues involved in these disputes, it is worth describing the standard argument for the BB hypothesis in a bit more detail. To simplify matters, for the most part we restrict attention to classical physics, since much of the controversy also imposes that restriction. The associated standard argument for the BB hypothesis starts by pointing out that we currently have two particular sets of information:
@@
[…]
@@
    The set $\cal D$ of all present observations and that we have concerning the physical universe. This includes all of our current records, recorded data, and memories.
[…]
Building on the widely accepted set of laws of modern classical and quantum physics, the fluctuation theorems of stochastic thermodynamics and open quantum thermodynamics have established that the evolution of entropy in all statistical physics systems, is a stochastic process Shiraishi (2023); Peliti and Pigolotti (2021); Seifert (2025). It is also widely accepted that (in non-cosmological contexts) this process:
  1. 1.
    is time translation invariant;
  2. 2.
    is symmetric under time-inversion;
  3. 3.
    has the property that if it is conditioned on a value of entropy $s_{0}$ at one particular time, $t_{0}$, and not conditioned on anything else, and if $s_{0}$ is sufficiently lower than the maximal value of entropy, then the expected value of entropy for times greater than $t_{0}$ are greater than $s_{0}$.
We call this set of three presumptions the entropy conjecture. We call any process satisfying the first two of these presumptions a Boltzmann process (whether or not it involves physical entropy).
@@
In addition to implying that the universe’s entropy evolves as a Boltzmann process, $\cal D$ includes the current value of the universe’s entropy, and in particular the fact that it is far lower than its maximum. (In other words, our current data tells us the current value of the universe’s entropy, albeit only with limited precision.) Crucially, the H theorem says that with extraordinarily high probability, entropy increases both into the future and into the past, from any such time at which entropy is known to have some specific low value. (Indeed, the entropy conjecture is time-symmetric about such a special time.) In other words, the most probable situation, given the current observations, is that we happen to be precisely at a special point in the dynamics of the universe’s entropy. In other words, the most probable situation is that we are just an entropy fluctuation, which is to say that we are a BB.
@@
This standard argument for the BB hypothesis initially arose in the context of de Sitter quantum cosmology Albrecht and Sorbo (2004); Linde (2007); Goheer et al. (2003); Dyson et al. (2002). If such a universe evolves into the future, it asymptotes to a vacuum state with a finite, non-zero temperature. This equilibrium state has fluctuations that act similarly to thermal fluctuations in a classical gas at equilibrium, which satisfy an entropy conjecture and corresponding recurrence theorems.
[…]
One may respond to these arguments that we also know the second law of thermodynamics, and from this we know that entropy must grow with $t$. This directly contradicts the BB hypothesis. However, our knowledge of the second law arose from consideration of our data records about the past: how do we know that these are data records of the past and aren’t themselves due to a fluctuation? A common response is to simply assume that entropy was in fact low in the past and has been growing since. But this amounts to a (very) improbable assumption, given the entropy conjecture and $\cal D$. Those data $\cal D$ are vastly more likely to be the result of a fluctuation than to be the result of low entropy in the distant past.
Round and round the arguments go. A common problem with all these arguments is that they are, at best semi-formal. The only way to resolve their apparent conflicts, is to use fully rigorous reasoning, grounded in stochastic process theory. That is what we do here.
In a nutshell, our analysis starts by recognizing that what Boltzmann proved (though he did not have the tools to formulate this way) is that the stochastic process of entropy (or in his formulation, the H quantity) is a time-stationary Markov process, symmetric under time inversion, which is biased towards high values.
Crucially, to compute a distribution over possible trajectories of a random variable that evolves according to any stochastic process, it is first necessary to say precisely what (if anything) that distribution is conditioned on. In the context of the entropy conjecture, that means specifying a (perhaps empty) set of pairs of a value of entropy, and an associated time when the universe had that value of entropy.
Any such specified conditioning events cannot arise from the entropy conjecture. In fact, it is not provided by physics. Rather such events must be determined using Bayesian reasoning, based on our current observational data and the prior we adopt over the possible laws of the universe.
As we show, this simple observation, formulated in a fully rigorous manner, disentangles all the ambiguities and circular reasoning that are rife in the literature on Boltzmann brains, and more generally in the literature on the past hypothesis (PH), and even in the literature on the second law. In particular, our analysis shows that the PH and the BB hypothesis are, formally speaking identical, in that both only condition on a subset of our current data, differing only in which such data they condition on. We also show that arguments about “unstable reasoning” prove nothing concerning the validity of the BB hypothesis, only about one particular argument that has been offered for that validity. We also formalize some of the properties of all circular reasoning, illustrating that they apply to common arguments for the second law.
We emphasize that we do not make any arguments for or against the BB hypothesis, the PH, or the second law. Indeed, as we prove, ultimately any such arguments must rely on choices about what data to condition the H theory on, and that choice cannot be provided by physics. Rather we provide a novel, fully formal framework for investigating those concepts and their very subtle relations.
@@
We also emphasize that we do not dispute the major role of cosmology in the second law of thermodynamics Rovelli (2019); Davies (1977); Scharnhorst and Aguirre (2023) or investigate precisely which models predict BB and which do not. Indeed, there are even major arguments that a framework for studying BB depending strictly on the entropy conjecture, without cosmology, is incomplete. It has long been known that the BB issue is highly sensitive to schema for assigning probabilities to values of the scale factor in FLRW scalar field cosmology models Page (2008); De Simone et al. (2010); Vilenkin (2007). We simply start from the working basis of pure statistical physics to unify and contextualize a recent genre of arguments in the literature.
We begin in the next section by summarizing one of the more nuanced, semi-formal counter-arguments that have been made to the BB hypothesis. A central concern with the counter-argument presented there is how we can reason from current data to conclusions about how experiments were initialized in the past, and from there to infer laws of the universe. The key conundrum we highlight is that this inference process itself relies on those laws being inferred, and so is circular. (Arguments in the literature that reach similar conclusions have characterized the standard argument for the BB hypothesis as “ unstable reasoning” Myrvold (2016); Carroll (2016) for some work related to this counter-argument.)
In the following section we begin by introducing the mathematical machinery we will need to disentangle the arguments concerning the BB hypothesis. Our first contribution is to introduce a restriction on sigma algebras that is necessary if (as in all analyses of of the PH, the second law, etc.) we use to investigate probability measures over possible “laws of physics”. We then show how to fully formalize the entropy conjecture as a Markov process; this is our second contribution. It builds crucially on the results in Scharnhorst et al. (2024).
Our third contribution is to use this formalization of the entropy conjecture to clarify what it cannot tell us. In contrast, our fourth contribution is to clarify what the entropy conjecture does in fact tell us concerning the BB hypothesis specifically. Next, our fifth contribution is to use our formalization of the entropy conjecture to analyze what arguments concerning unstable reasoning can and cannot tell us about the BB hypothesis. After this we present our sixth contribution, which is a variant of the BB hypothesis we call the “1000 CE hypothesis”. As we show, formally speaking, the 1000 CE hypothesis lies exactly between the standard BB hypothesis and the PH. This illustrates the underlying identity between standard BB hypothesis and the PH.
We end with a discussion, and then present some of the myriad future areas of research that our investigation suggests.
II An objection to the standard argument for Boltzmann Brains
[…]
To recap from the introduction, the conditional probability $P({\cal B}\,|\,{\cal L},{\cal D})$ of the BB hypothesis, given the laws of physics and the data observed in the present, is near $1$. So in terms of our shorthand, that standard argument says that
@@ II An objection to the standard argument for Boltzmann Brains
[…]
@@
[…]
$P({\cal B}\,|\,{\cal{D}},{\cal R})\simeq P({\cal B}\,|\,{\cal{D}},{\cal R},{\cal L})P({\cal L}\,|\,{\cal{D}},{\cal R}),$ (4)
Even though the second term on the RHS, $P({\cal L}\,|\,{\cal{D}},{\cal R})$, is close to $1$ (by hypothesis), the first term on the RHS, $P({\cal B}\,|\,{\cal{D}},{\cal R},{\cal L})$, is close to 0. Therefore $P({\cal B}\,|\,{\cal{D}},{\cal R})$ is close to 0. Therefore, returning to our simplified notation,
@@
[…]
III Relation between the entropy conjecture, Boltzmann Brains, and the second law
The discussion in the preceding sections may leave a sense of confusion. Many parts of the arguments recounted above involving the BB hypothesis might seem to involve circular reasoning. Such reasoning can only (!) establish that our priors are inconsistent. However, the fact that an argument engages in circular reasoning doesn’t tell us anything about how valid its conclusions are, just that the priors it relies on don’t themselves establish those conclusions. (See Appendix B for a Bayesian definition “circular reasoning”, and an analysis of what it does and does not imply.)